OpenStudy (anonymous):

Will give fan & medal!!! Help please :) Which shows the expression below simplified? 0.036 ÷ (4 × 10-2) A)-0.4 × 10-1 B)9 × 10-1 C)9 × 100 D)9 × 10-2

OpenStudy (paxpolaris):

\[0.036\over4\times 10^{-2}\]

OpenStudy (paxpolaris):

\[={36\times 10^{-3}\over4\times 10^{-2}}\]

OpenStudy (paxpolaris):

\[\frac{36}4 \times 10^{-3-\left( -2 \right)}\]

OpenStudy (paxpolaris):

\[9\times 10^{-3+2}\]

OpenStudy (anonymous):

@PaxPolaris Thanks so much :)

OpenStudy (paxpolaris):

np :)